OpenStudy (anonymous):

Which set of numbers is an arithmetic sequence? A. 1, 5, 8, 13, ... B. 2, 4, 8, 32, ... C. 25, 21, 17, 13, ... D. 81, 27, 9, 3, ...

OpenStudy (dayakar):

i will give u an example 1,3,5,7,...

OpenStudy (dayakar):

check each one u can easily identify

jhonyy9 (jhonyy9):

for B. 2,4,8,32,... 2*4=8 4*8=32 8*32=?